OpenStudy (anonymous):

Write an equation for: Twelve times h equals one hundred and twenty minus thirty-six. State whether the equation is true, false, or an open sentence. Explain.

OpenStudy (anonymous):

times means multiply.. equals means = minus means - So: 12*h = 120 - 36..

OpenStudy (anonymous):

12(h) = 120 - 36

OpenStudy (anonymous):

i have the equation i dont understand the second part

OpenStudy (anonymous):

True

OpenStudy (anonymous):

Last one..

OpenStudy (anonymous):

Why it is true?/ Do you know the values of h ??

OpenStudy (anonymous):

its false ?

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

Neither it is false: for h = 7 you have equation equal.. It is the last one @skittles25

OpenStudy (anonymous):

so its an open sentence

OpenStudy (anonymous):

Yes ^

OpenStudy (anonymous):

Yes it is...

OpenStudy (anonymous):

why? i dont get it

OpenStudy (anonymous):

See it is not true because you don't know the values of h.. See, If h = 1 then tell me is the equation equal?/ LHS = RHS ??

OpenStudy (anonymous):

what wow im even more lost i just dont get why its an open sentnece

OpenStudy (anonymous):

See for equation to be equal, LHS should be equal to RHS.. 12*h = 120 - 36 Firstly tell me what is 120 - 36 ??

OpenStudy (anonymous):

Can you guys help me with this?Question 5 (Essay Worth 4 points) Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.

OpenStudy (anonymous):

Tell me..

OpenStudy (anonymous):

im so confused i just need and explaination for why its an open question

OpenStudy (anonymous):

@waterineyes im so confused i just need an explaination as to why its and open sentence